OpenStudy (anonymous):

solve 5x-49≥6-8(2x+7) please show work

OpenStudy (anonymous):

\[5x-49\ge6-16x-56\] \[5x-49\ge-16x-50\] \[5x+16x \ge-50+49\] \[21x \ge-1\] \[x \ge-1/21\]